OpenStudy (jocelynstalnaker):

Help? A quality control inspector is in charge of inspecting drill bits that are produced on an assembly line. The diameter of a drill bit can be modeled by the equation |x − 0.525| = 0.01. Show your work in solving the equation to determine the minimum and maximum diameters of the drill bits.

OpenStudy (jocelynstalnaker):

@Loser66 Its from my last module and I never understood it.

OpenStudy (loser66):

neither I. @zepdrix

OpenStudy (jocelynstalnaker):

Haha ok.

zepdrix (zepdrix):

\[\Large\rm \left|x-0.525\right|=0.01\]One way we can rewrite the absolute bars is like this:\[\Large\rm \pm\left(x-0.525\right)=0.01\]

OpenStudy (perl):